Filer Manufacturing has 4,916,190 shares of common stock outstanding. The current share price is $52.39, and the book value per share is $3.65. Filer Manufacturing also has two bond issues outstanding. The first bond issue has a face value of $62,308,724, has a 0.05 coupon, matures in 10 years and sells for 83 percent of par. The second issue has a face value of $61,013,614, has a 0.06 coupon, matures in 20 years, and sells for 92 percent of par.

The most recent dividend was $2.13 and the dividend growth rate is 0.06. Assume that the overall cost of debt is the weighted average of that implied by the two outstanding debt issues. Both bonds make semiannual payments. The tax rate is 0.32.

What is Filer's aftertax cost of debt? Enter the answer with 4 decimals (e.g. 0.2345)
Answer #1
Cost of debt
Bond1
                  K = Nx2
Bond Price =∑ [(Semi Annual Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                  K =10x2
51716240.92 =∑ [(5*62308724/200)/(1 + YTM/200)^k]     +   62308724/(1 + YTM/200)^10x2
                   k=1
YTM1 = 7.440163436
Bond2
                  K = Nx2
Bond Price   =∑ [(Semi Annual Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                   K =20x2
56132524.88 =∑ [(6*61013614/200)/(1 + YTM/200)^k]     +   61013614/(1 + YTM/200)^20x2
                    k=1
YTM2 = 6.73
Firm cost of debt=YTM1*(MV bond1)/(MV bond1+MV bond2)+YTM2*(MV bond2)/(MV bond1+MV bond2)
Firm cost of debt=7.440163436*(51716240.92)/(51716240.92+56132524.88)+6.73*(51716240.92)/(51716240.92+56132524.88)
Firm cost of debt=7.07%
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 7.07*(1-0.32)
= 0.4808